OpenStudy (anonymous):

Which is an example of an algebraic expression? (1 point) 4(3 + 8) 182 3 – a 21 • 4 2. Which is an algebraic expression for 5 more than z? (1 point) 5z z + 5 5z + 5 5 – z 3. Which word phrase can be used to represent the algebraic expression 4(21 + n)? (1 point) 4 plus the sum of 21 and a number n 4 times the product of 21 and a number n 4 times the sum of 21 and a number n 4 less than the sum of 21 and a number n 4. Which is an algebraic expression for the word phrase "8 less than the product of 12 an 2 ye

OpenStudy (anonymous):

1= 3-a 2= z+5 3= 4 times the sum of 21 and a number n 4= 8< 12*2